A person consumes 70 grams of fat, 40 grams of carb, 25 grams of protein, 15...

  1. A person consumes 70 grams of fat, 40 grams of carb, 25 grams of protein, 15 milligrams of thiamin, and 250 milligrams of calcium. How much energy did that person consume?
  1. A woman consumes 400 grams of carb, 50 grams of protein, 65 grams of fat in one day.

     How many total calories did she consume, and how many and what percentage of calories were from carbohydrate ,protein and fat?
Answer #1

Solution:
Carbohydrate = 4 calories per gram
Protein = 4 calories per gram
Fat = 9 calories per gram

Thaimin and calcium do not produce any calorie because vitamins and minerals do not produce calories and all the calories used by body is produce through fat, carbhohydrate and protein.

Person consumes 70 grams of fat, 40 grams of carb, 25 grams of protein,hence calories consumes by Man
= 70 × 9 + 40 × 4 + 25 × 4 = 630 + 160 + 100 = 890

Woman consumes 400 grams of carb, 50 grams of protein, 65 grams of fat in one day.
Calories consumes by women
= 400 × 4 + 50 × 4 + 65 × 9
= 1600 + 200 + 585
= 2385

(By woman)100% calories are produce from fats,carbohydrates and proteins.
Percentage of carb = 1600 × 100/2385 = 67.08%

Percentage of protein = 200 × 100/ 2385 = 8.38 %


Percentage of fat = 585 × 100/2385 = 24.52 %